600 invested for 3 years at 25%
What is the interest per year

Answer by JulietG(1812)

I'd love to see the prospectus for an investment that pays 25%!
This assumes that it is simple interest, paid yearly.
First year, 600 * 0.25 = 150 interest
Second year (600+150) * 0.25 = 187.50 interest
Third year (600+150+187.50) * 0.25 = 234.38 interest
.
At the end of the third year, you'll have
600 (initial) + 150 (yr 1 interest) + 287.50 (yr 2 interest) + 187.50 (yr 3 interest)